# Towards Geometric-Photometric Joint Alignment for Facial Mesh Registration

### v1

@@ Abstract
This paper presents a Geometric-Photometric Joint Alignment(GPJA) method, for accurately aligning human expressions by combining geometry and photometric information. Common practices for registering human heads typically involve aligning landmarks with facial template meshes using geometry processing approaches, but often overlook photometric consistency. GPJA overcomes this limitation by leveraging differentiable rendering to align vertices with target expressions, achieving joint alignment in geometry and photometric appearances automatically, without the need for semantic annotation or aligned meshes for training. It features a holistic rendering alignment strategy and a multiscale regularized optimization for robust and fast convergence. The method utilizes derivatives at vertex positions for supervision and employs a gradient-based algorithm which guarantees smoothness and avoids topological defects during the geometry evolution. Experimental results demonstrate faithful alignment under various expressions, surpassing the conventional ICP-based methods and the state-of-the-art deep learning based method. In practical, our method enhances the efficiency of obtaining topology-consistent face models from multi-view stereo facial scanning.
@@ 1 Introduction
Nowadays, professional studios in industry and academia commonly use synchronized multi-view stereo setups for facial scanning [17, 36, 44], ensuring high-fidelity results in controlled settings. These setups aim to generate topology-consistent meshes for different subjects with various facial expressions. Typically, conventional pipelines [13] involve constructing raw scans from multi-view images, followed by manual processes like marker point tracking, clean-up, or key-framing [4], which is labor-intensive and time-consuming, limiting its application in film, gaming, AR/VR industry. To fulfil automatic registration, geometry based methods have been widely employed [1, 21]. However, these methods primarily focus on geometric alignment, neglecting the photometric consistency. To remedy this issue, this paper aims to achieve a joint alignment in terms of geometry and photometric appearances. To this end, two challenges need to be addressed.
The first challenge is to establish a proper deformation field to guide the alignment process, especially for the challenging areas such as mouths and eyes. Previous attempts have been made to construct correspondences by landmarks or optical flow to aid the photomeric alignment [13, 1]. However, the offset vectors obtained through theses approaches in the 2D image space are insufficient in guiding deformation in 3D world. Gafni et al. [14] employ implicit volumetric representation to combine shape and appearance recovery for realistic rendering, which lacks explicit geometry constraints. In response to this challenge, we propose a differentiable rendering  [20, 31] based registration framework to generate topology-consistent facial meshes from multi-view images. In particular, our approach includes a Holistic Rendering Alignment(HRA) which incorporates constraints from color, depth and surface normals, facilitating alignment through automatic differentiation without explicit correspondence computation.
@@
Figure 1: Given multi-view images and (a) the textured template mesh, we propose a novel method GPJA based on differentiable rendering to achieve geometric and photometric alignment jointly for facial meshes. (b) The aligned meshes are rendered with the shared texture map as the template in (a). (c) The zoomed renderings of eyes and mouths demonstrate photometric alignment with the reference images.
@@
The second challenge for facial mesh registration lies in generating faithful output meshes while retaining the topological structure. Aligning facial expressions involves large-step geometry deformation, and thus is prone to topological artifacts. In light of this problem, we resort to a Multiscale Regularized Optimization that combines a modified gradient descent algorithm [30] with coarse-to-fine remeshing scheme. Starting with the coarsest template, the mesh is tessellated periodically while updating the vertices with an efficient regularized geometry optimization for the contraints collected from holistic rendering alignment. Our multiscale regularized optimization ensures smoothness and fast convergence without requiring a specific amount of training data.
@@
We validate our method through experiments on six subjects from the Face-Scape [43] dataset, covering diverse facial expressions. The joint alignment is examined by both geometric and image metrics, demonstrating the effectiveness of our approach. The aligned meshes produced by our method are of high quality, free from topological errors, and accurately warped even in challenging regions like mouths and eyes.
@@
  • A new method named GPJA achieving joint alignment in geometry and photometric appearances for facial meshes, without any semantic annotation like facial landmarks.
  • A holistic rendering alignment module based on differentiable rendering that effectively generates the deformation field for joint alignment.
  • A multiscale regularized optimization that produces high-quality aligned meshes using an efficient gradient-based algorithm.
@@ 2 Related Work
Our research centers on the registration of facial meshes for topology-consistent geometry. This section provides a literature review relevant to our study.
@@
Geometry Processing Methods. Non-rigid registration is a well-established technique in geometry processing for warping a template mesh to raw scans [40, 37, 3]. The strong geometric prior of the template model enhances local shape matching and enables large-step deformation, making it suitable for facial mesh processing. The Iterative Closest Point(ICP) algorithm is a commonly used framework [1, 37, 22] for registration. With a template based on 3D Morphable Models [43, 16] as initialization, ICP minimizes the error between landmarks on the template and the scans, resulting in a rough alignment. Then, a fine-tuning stage involves searching for valid correspondences in the spatial neighborhood, and warping the template leveraging data fidelity and smoothness terms. Previous work has explored regularization terms [42, 32] and correspondences [41] in ICP-based algorithms. However, these methods [43, 9, 16] have limitations in achieving photometric consistency as they rely solely on geometric information.
Figure 2: Illustration of our proposed GPJA. With the provided textured template, scans and camera poses, the contraints from holistic rendering alignment $\mathcal{L}$ back-propagates derivatives at each iteration to guide the warping of the template. The regularized optimization is built on a multiscale scheme with periodic tessellation, and the vertices are updated with an efficient modified gradient descent algorithm.
@@
To address photometric alignment, researchers have explored incorporating optical flow [13, 33, 8] into the fine-tuning stage of facial mesh registration. However, optical flow alone is inadequate for handling significant differences between the source template and target scans, as well as occlusion changes around the eyes and mouth. Another challenge in facial mesh registration is maintaining smooth contours in facial features [5], which is difficult due to occlusions and color changes. Previous approaches have used user-guided methods [12] or contour extraction [15] to address these challenges, but they either involve lengthy workflows or specific treatments.
@@
Deep Learning Based Methods. Recent advancements in deep learning approaches for generating topology-consistent meshes include ToFu [23], which predicts probabilistic distributions of individual vertices of a template face mesh to reconstruct registered face geometry. TEMPEH [6] enhances ToFu with a transformer-based architecture, while NPHM [16] models head geometry using a neural field representation that parameterizes shape and expressions in disentangled latent spaces. While these methods prioritize geometric alignment, they do not guarantee rigid photometric consistency. ReFA [25] introduces a recurrent network operating in the texture space for predicting positions and texture maps. Although these deep learning methods represent progress in facial mesh registration, they require a substantial amount of registered data processed with classical ICP-based algorithms.
@@
In addition to mesh-based representations, implicit volumetric representations have gained popularity in reconstruction. Several studies have extended NeRF [27] for dynamic face reconstruction [34, 45, 2]. The pipeline typically begins with explicit parametric models, followed by estimating a deformation field. Finally, a volumetric renderer is used to generate densities and colors. However, NeRF-style methods lack supervision for aligning explicit-represented geometry and generally do not produce production-ready geometries despite decent rendering results [25].
@@ 3 Preliminaries
Before discussing the details of our methodology, we first give a brief introduction on the differentiable rendering to ease the understanding of our method.
[…]
$\boldsymbol{I}(\boldsymbol{p}(x,y))=F(\boldsymbol{x};\Theta),$ (1)
where the function $F(\cdot)$ represents the rendering process, encompassing various computations such as shading, interpolation, projection, and antialiasing. The output of this function can be RGB colors, normals, depths, or label images etc.. In our scenario, the parameter to be optimized is the positions of mesh vertices denoted by $\boldsymbol{x}\in\mathbb{R}^{n\times 3}$. $\Theta$ symbolizes a set of scene parameters known in advance, including camera poses, lighting, texture color, and other relevant factors.
Differentiable rendering augments renderers by additionally providing derivatives with respect to certain scene parameters, which is an emerging tool for inverse problems [31, 28]. The objective of inverse rendering is to recover some specific scene parameters through gradient-based optimization on a scalar loss function $\mathcal{L}$ which is usually defined as the sum of pixel-wise differences between the rendered images and reference images $\boldsymbol{I}^{\textit{{r}}}\in\mathbb{R}^{w\times h\times c}$ across $v$ views. In this work, we adopt the $L_{1}$-norm for loss functions.
$\mathcal{L}\left(\boldsymbol{x}\mid\Theta,\boldsymbol{I}^{\textit{{r}}}\right)=\sum_{j}^{v}\left|F(\boldsymbol{x}\mid\Theta_{j})-\boldsymbol{I}_{j}^{\textit{{r}}}\right|.$ (2)
In our setting of joint alignment registration, we use the derivative $\frac{\partial\mathcal{L}}{\partial\boldsymbol{x}}$ to guide the template mesh in fitting the target expressions while maintaining topological consistency.
@@ 4 Joint Alignment of Facial Meshes
Figure 2 depicts the pipeline of our method. Starting with the provided textured template mesh and raw scans, each iteration first computes deformation from the holistic rendering alignment and then optimize the vertices via the multiscale regularized optimization.
@@ 4.1 Holistic Rendering Alignment
Motivation. Using the provided template mesh and raw scans, non-rigid 3D registration calculates the deformation parameterized on the vertices of the template mesh to align with the target scan. While the geometric alignment has been well studied [10], previous photometric alignment attempts mainly learn 2D-constrained deformations [33, 13] expressed in the image space, which are prone to false correspondences and ambiguities when elevating to the 3D space. On the other hand, we note that differentiable rendering is capable of providing automatically computed derivatives $\frac{\partial\mathcal{L}}{\partial\boldsymbol{x}}$ as guidance to deform the template mesh ${\boldsymbol{T}}$ into joint alignment, avoiding correspondence errors from semantic facial features [1] or offset vectors by optical flow [33]. Furthermore, differentiable rendering retrieves the deformation directly in the 3D space, addressing occlusion issues in eye and mouth regions.
@@
$\mathcal{L}=\mathcal{L}_{color}+\mathcal{L}_{depth}+\mathcal{L}_{normal}.$ (3)
Color Constraint. $\mathcal{L}_{color}$ seeks to impose constraints for photometric alignment by comparing the rendered image with the observed multi-view color images. However, the deformation on the inner lip contours is occasionally disturbed by occlusion changes. To address this issue, we deliberately exclude the interior mouth from the color constraint by a masking strategy.
@@
$\begin{split}\mathcal{L}_{color}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{S}\left(\boldsymbol{x}\mid\boldsymbol{P}_{j},\boldsymbol{S},\boldsymbol{C}_{\boldsymbol{T}}\right)-\boldsymbol{I}_{j}^{\textit{{r}}}\right)\odot\\ F_{S}\left(\boldsymbol{x}\mid\boldsymbol{P}_{j},\boldsymbol{B}\right)|,\end{split}$ (4)
where the shading function $F_{S}:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{w\times h\times 3}$ is defined as rendering diffuse objects with the lighting estimated from the capture setup in spherical harmonics forms $\boldsymbol{S}\in\mathcal{\mathbb{R}}^{9\times 3}$ [35], $\boldsymbol{P}_{j}\in\mathbb{PL}\left(3\right)$ remarks projective matrix of the $j$-th camera, and $\odot$ denotes element-wise multiplication. By synthesizing a binary image using $F_{S}(\boldsymbol{x}|\boldsymbol{P}_{j},\boldsymbol{B})$, where the interior mouth is assigned zeros and the rest with ones, Eq. 4 can mask out the interior mouth for color constraint. Experimental results confirm the effectiveness of this mask operation.
Figure 3: Illustration of the mask operation. Using the same parameterization as (a) the texture map, the interior mouth is manually masked out, resulting in (b) a binary mask image. By the shading operation $F_{S}(\boldsymbol{x}|\boldsymbol{P}_{j},\boldsymbol{B})$, (c) the mouth socket is mask out for color constraint.
@@
$\mathcal{L}_{depth}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{D}\left(\boldsymbol{x}\mid\boldsymbol{P}_{j}\right)-F_{D}\left(\boldsymbol{\tilde{x}}\mid\boldsymbol{P}_{j}\right)\right)|,$ (5)
@@
$\mathcal{L}_{normal}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{N}\left(\boldsymbol{x}\mid\boldsymbol{P}_{j}\right)-F_{N}\left(\boldsymbol{\tilde{x}}\mid\boldsymbol{P}_{j}\right)\right)|,$ (6)
@@
The holistic rendering alignment benefits from differentiable rendering to obtain derivatives, replacing the explicitly computed correspondences from previous approaches [41, 10]. By incorporating multiple cues, holistic rendering alignment ensures joint alignment of geometry and photometric appearances.

### v2

@@ Abstract
This paper presents a Geometric-Photometric Joint Alignment (GPJA) method, which aligns discrete human expressions at pixel-level accuracy by combining geometric and photometric information. Common practices for registering human heads typically involve aligning landmarks with facial template meshes using geometry processing approaches, but often overlook dense pixel-level photometric consistency. This oversight leads to inconsistent texture parametrization across different expressions, hindering the creation of topologically consistent head meshes widely used in movies and games. GPJA overcomes this limitation by leveraging differentiable rendering to align vertices with target expressions, achieving joint alignment in both geometry and photometric appearances automatically, without requiring semantic annotation or pre-aligned meshes for training. It features a holistic rendering alignment mechanism and a multiscale regularized optimization for robust convergence on large deformation. The method utilizes derivatives at vertex positions for supervision and employs a gradient-based algorithm which guarantees smoothness and avoids topological artifacts during the geometry evolution. Experimental results demonstrate faithful alignment under various expressions, surpassing the conventional non-rigid ICP-based methods and the state-of-the-art deep learning based method. In practical, our method generates meshes of the same subject across diverse expressions, all with the same texture parametrization. This consistency benefits face animation, re-parametrization, and other batch operations for face modeling and applications with enhanced efficiency.
keywords:
Geometry registration , Facial performance capture , Face modeling
journal: Computers & Graphics
[…]
@@ 1 Introduction
Figure 1: Given multiview images and (a) the textured template mesh, we propose a novel method GPJA based on differentiable rendering to achieve geometric and photometric alignment jointly for facial meshes. (b) The aligned meshes are rendered with the shared texture map as the template in (a). (c) The zoomed renderings (yellow boxes) of eyes and mouths demonstrate photometric alignment with the reference images (black boxes).
Nowadays, professional studios in industry and academia commonly use synchronized multiview stereo setups for facial scanning [1, 2, 3], ensuring high-fidelity results in controlled settings. These setups aim to generate topology-consistent meshes for different subjects with various facial expressions. Typically, conventional pipelines [4] involve constructing raw scans from multiview images, followed by manual processes like marker point tracking, clean-up, or key-framing [5], which is labor-intensive and time-consuming, limiting its application in film, gaming, AR/VR industry. To fulfill automatic registration, geometry-based methods have been widely employed [6, 7, 8, 9, 10]. However, these methods primarily focus on geometric alignment but failing to ensure photometric consistency at dense pixel-level. To remedy this issue, this paper aims to achieve a joint alignment in terms of geometry and photometric appearances. To this end, two challenges need to be addressed.
The first challenge is to establish a proper deformation field to guide the alignment process, especially for the challenging areas such as mouths and eyes. Previous attempts have been made to construct correspondences by landmarks or optical flow to aid the photometric alignment [4, 6]. However, the offset vectors obtained through these methods often introduce errors. Moreover, since they are extracted from 2D images, these vectors are insufficient for guiding deformation of 3D geometry.  [11] employed implicit volumetric representation to combine shape and appearance recovery for realistic rendering, which lacks explicit geometry constraints. In response to the challenge, we propose a differentiable rendering [12, 13] based registration framework to generate topology-consistent facial meshes from multiview images. In particular, our approach includes a Holistic Rendering Alignment (HRA) which incorporates constraints from color, depth and surface normals, facilitating alignment through automatic differentiation without explicit correspondence computation.
@@
The second challenge for facial mesh registration is generating faithful output meshes while preserving the topological structure. Aligning discrete facial expressions involves large-step geometry deformation, which is susceptible to topological artifacts. Previous works addressed this with specialized constraints [8, 10], but at the cost of a complex pipeline. Other approaches [14, 15, 7, 16] utilize parametric models to fit the facial images for mesh creation. Although these methods are highly efficient, they struggle with diverse identities and complex expressions due to their limited expressive capabilities. To overcome this, we resort to a multiscale regularized optimization that combines a modified gradient descent algorithm [17], with coarse-to-fine remeshing scheme. Starting with the coarsest template, the mesh is tessellated periodically while updating the vertices with a robust regularized geometry optimization for the constraints collected from HRA. Our multiscale regularized approach is simple to implement and ensures smoothness and robust convergence, without the need for extensive training data.
@@
We validate our method through experiments on seven subjects from the FaceScape [18] dataset, covering diverse facial expressions, as shown in Fig. 1. The joint alignment is examined by both geometric and image metrics, demonstrating the effectiveness of our approach. The aligned meshes produced by our method are of high quality, free from topological errors, and accurately warped even in challenging regions like mouths and eyes.
@@
  • 1.
    A simple and novel method named GPJA achieving joint alignment in geometry and photometric appearances at dense pixel-level for facial meshes.
  • 2.
    A holistic rendering alignment mechanism based on differentiable rendering that effectively generates the deformation field for joint alignment, without any semantic annotation.
  • 3.
    A multiscale regularized optimization strategy that ensures robust convergence and produces high-quality, topologically consistent aligned meshes.
@@ 2 Related Work
Our research focuses on the registration of facial meshes for topology-consistent geometry on discrete expressions. This section provides a literature review relevant to our study.
@@
Geometry Processing Methods. Non-rigid registration is a well established technique in geometry processing for warping a template mesh to raw scans [19, 20, 21, 22]. The Iterative Closest Point (ICP) algorithm is a commonly used framework [6, 20, 23] for registration. With a template based on 3D Morphable Models (3DMMs) of strong geometric priors [6, 24, 7, 8, 9, 18, 25, 10] as initialization, ICP minimizes the error between landmarks on the template and the scans, resulting in a rough alignment. Then, a fine-tuning stage involves searching for valid correspondences in the spatial neighborhood, and warping the template leveraging data fidelity and smoothness terms. Previous works have explored regularization terms [26, 27] and correspondences [28, 29] in ICP-based algorithms. In addition to registration, there are also face tracking methods [30, 31] that rely on coefficient regression of rigged blendshapes, which are primarily used for expression transfer rather than high-fidelity reconstruction. Overall, these methods are limited in achieving pixel-level photometric consistency, where our method is successfully capable of.
To achieve photometric alignment, industry-standard methods often involve re-topologizing frame-by-frame using the professional software like Wrap4D [32], which requires significant time and resources [33]. For automatic pipelines, researchers have explored incorporating optical flow [34, 35, 36] into the fine-tuning stage of facial mesh registration. However, optical flow alone is inadequate for handling significant differences between the source template and target scans, as well as occlusion changes around the eyes and mouth [37]. It is even worse for discrete expression scenarios due to the lack of temporal coherence and large deformation. In contrast, our method effectively handles significant deformation, visibility changes and asymmetry.
@@
Another challenge in facial mesh registration is maintaining smooth contours in facial features [38], which is difficult due to occlusions and color changes. Previous approaches have used user-guided methods [33] or contour extraction [37] to address these challenges, but they either involve lengthy workflows or specific treatments, which hinder efficient topology-consistent mesh creation. On the other hand, our method achieves pixel-level alignment without any semantic annotation.
@@
Deep Learning Based Methods. Recent advancements in deep learning approaches for generating topology-consistent meshes include ToFu [39], which predicts probabilistic distributions of vertices on the face mesh to reconstruct registered face geometry. TEMPEH [40] enhances ToFu with a transformer-based architecture, while NPHM [41] models head geometry using a neural field representation that parametrizes shape and expressions in disentangled latent spaces. Although these methods prioritize geometric alignment, they do not guarantee rigorous photometric consistency. ReFA [42] introduces a recurrent network operating in the texture space for predicting positions and texture maps. Although these deep learning methods represent progress in facial mesh registration, they require a substantial amount of registered data processed with classical ICP-based algorithms. Besides, out-of-domain expressions may be limited by insufficient training data. In our work, no additional training data is needed except for a semi-automatic created template mesh with textures.
@@
In addition to mesh-based representations, implicit volumetric representations have gained popularity in reconstruction. Several studies have extended NeRF [43] and 3D Gaussian Splatting [44] for dynamic face reconstruction [45, 46, 47, 48, 49]. The pipeline typically begins with explicit parametric models, followed by estimating a deformation field represented by multi-layer perceptrons. Finally, a volumetric renderer is used to generate densities and colors. However, volumetric rendering-based methods lack supervision for aligning mesh-represented geometry and generally do not produce production-ready geometries despite decent rendering results [42].
@@ 3 Preliminaries
Before discussing the details of our methodology, we first introduce differentiable rendering as background information to provide context for our approach.
Given a 3D scene containing mesh-based geometries, lights, materials, textures, cameras etc., a renderer synthesizing a 2D image P of each screen pixel (x,y) can be formulated as:
$\boldsymbol{P}(x,y)=F(\boldsymbol{x}|\Theta),$ (1)
where the function $F(\cdot)$ represents the rendering process, encompassing various computations such as shading, interpolation, projection, and anti-aliasing. The output P of this function can be RGB colors, normals, depths, or label images. In our scenario, the parameter to be optimized is the positions of mesh vertices denoted by $\boldsymbol{x}\in\mathbb{R}^{n\times 3}$. $\Theta$ symbolizes a set of scene parameters known in advance, including camera poses, lighting, texture color, and other relevant factors.
Differentiable rendering augments renderers by providing additional derivatives with respect to certain scene parameters, which is a valuable tool for inverse problems [13, 50]. The objective of inverse rendering is to recover some specific scene parameters through gradient-based optimization on a scalar loss function $\mathcal{L}$ which is usually defined as the sum of pixel-wise differences between the rendered images $\boldsymbol{P}_{j}$ and the reference image $\boldsymbol{P}_{j}^{\textit{{r}}}\in\mathbb{R}^{w\times h\times c}$ at the j-th camera pose across $v$ views. In this work, we adopt the $L_{1}$-norm for loss functions.
$\mathcal{L}\left(\boldsymbol{x}\mid\Theta,\boldsymbol{P}^{\textit{{r}}}\right)=\sum_{j}^{v}\left|F(\boldsymbol{x}\mid\Theta_{j})-\boldsymbol{P}_{j}^{\textit{{r}}}\right|.$ (2)
In our joint alignment registration setting, the derivative $\frac{\partial\mathcal{L}}{\partial\boldsymbol{x}}$ obtained through differentiable rendering plays a key role, as it guides the template mesh to fit the target expressions. This process avoids the explicit computation of correspondences, such as optical flow, facial landmarks, or other semantic information.
@@ 4 Joint Alignment of Facial Meshes
Figure 2: Illustration of our proposed GPJA. With the provided textured template, scans and camera poses, the constraints from HRA $\mathcal{L}$ back-propagates derivatives at each iteration to guide the warping of the template. The regularized optimization is built on a multiscale scheme with periodic tessellation, and the vertices are updated with a robust modified gradient descent algorithm.
Our work addresses the challenge of registering the same subject across diverse facial expressions. Fig. 2 depicts the pipeline of our method. Starting with a textured template mesh and raw scans, each iteration computes deformation from HRA and optimizes the vertices through multiscale regularized optimization. As a result, the outputs of GPJA share a joint alignment across all expressions within the subject.
@@ 4.1 Holistic Rendering Alignment
Motivation. While geometric alignment has been well studied, previous photometric alignment approaches mainly consider sparse facial landmarks, resulting in two main limitations: coarse alignment lacking fine details and unreliability under occlusions and extreme expressions, as shown in Fig. 3(a). Other photometric algorithms [35, 4] attempt to learn deformation from optical flow obtained in the 2D images, but are prone to false correspondences as shown in Fig. 3(b) ,and suffer from ambiguities when elevating to the 3D space.
(a) Common errors with landmarks
(b) Common errors with optical flow
Figure 3: Errors are indicated with arrows: (a) Landmark [51] errors under asymmetry, occlusion, and extreme expressions, and (b) Optical flow [52] inaccuracies due to significant visibility changes.
Instead of making remedies for landmarks or optical flows, we leverage differentiable rendering, which automatically computes derivatives $\frac{\partial\mathcal{L}}{\partial\boldsymbol{x}}$ to guide the deformation of the template mesh ${\boldsymbol{T}}$ into both geometric and photometric alignment. This approach eliminates correspondence errors from facial landmarks [6] or optical flow offset vectors [35]. Furthermore, differentiable rendering operates directly in 3D space, effectively handling occlusions in the eye and mouth regions.
@@
$\mathcal{L}_{HRA}=\mathcal{L}_{color}+\mathcal{L}_{depth}+\mathcal{L}_{normal}.$ (3)
Color Constraint. $\mathcal{L}_{color}$ seeks to impose constraints for photometric alignment by comparing the rendered image with the observed multiview color images. In practice, we found that the deformation around the inner lip is sometimes affected by occlusion changes, resulting in inaccurate lip contours. To address this issue, we deliberately exclude the interior mouth from the color constraint by a masking strategy.
@@
Figure 4: Illustration of the masking strategy. Using the same parametrization as (a) the texture map, the interior mouth is manually masked out, resulting in (b) a binary mask image. By the shading operation $F_{S}(\boldsymbol{x}|\boldsymbol{P}_{j},\boldsymbol{B})$, (c) the mouth socket is mask out from the color constraint.
@@
$\begin{split}\mathcal{L}_{color}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{S}\left(\boldsymbol{x}\mid\boldsymbol{\Pi}_{j},\boldsymbol{S},\boldsymbol{C}_{\boldsymbol{T}}\right)-\boldsymbol{P}_{j}^{\textit{{r}}}\right)\odot F_{S}\left(\boldsymbol{x}\mid\boldsymbol{\Pi}_{j},\boldsymbol{B}\right)|,\end{split}$ (4)
where the shading function $F_{S}:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{w\times h\times 3}$ is defined as rendering diffuse objects with the lighting estimated from the capture setup in spherical harmonics forms $\boldsymbol{S}\in\mathcal{\mathbb{R}}^{9\times 3}$ [53], $\boldsymbol{\Pi}_{j}\in\mathbb{PL}\left(3\right)$ remarks projective matrix of the $j$-th camera, and $\odot$ denotes element-wise multiplication. By synthesizing a binary image using $F_{S}(\boldsymbol{x}|\boldsymbol{\Pi}_{j},\boldsymbol{B})$, where the interior mouth is assigned zeros and the rest with ones, Equation 4 can mask out the interior mouth for color constraint for faithful lip contours. Experimental results confirm the effectiveness of this mask operation.
@@
$\mathcal{L}_{depth}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{D}\left(\boldsymbol{x}\mid\boldsymbol{\Pi}_{j}\right)-F_{D}\left(\boldsymbol{\tilde{x}}\mid\boldsymbol{\Pi}_{j}\right)\right)|,$ (5)
@@
$\mathcal{L}_{normal}\left(\boldsymbol{x}\right)=\sum_{j}^{v}|\left(F_{N}\left(\boldsymbol{x}\mid\boldsymbol{\Pi}_{j}\right)-F_{N}\left(\boldsymbol{\tilde{x}}\mid\boldsymbol{\Pi}_{j}\right)\right)|,$ (6)
@@
HRA benefits from differentiable rendering to obtain derivatives, replacing the explicitly computed correspondences from previous approaches [28, 56]. By incorporating multiple cues, HRA ensures joint alignment of geometry and photometric appearances.